(USE EXCEL) Grass Whacker’s is considering whether or not to refund a $90 million, 6 percent coupon, 30 year bond issue that was sold 8 years ago. It is amortizing $1.8 million of flotation costs over the issue’s 30-year life. A new 22-year issue would carry an interest rate of 5.35 percent. A call premium of 5 percent would be required to retire the old bonds and flotation costs on the new issue would be $1.35 million (also to be amortized). Grass Whacker’s marginal tax rate is 34 percent. The new bonds would be issued 1 month before the old bonds are called, with the proceeds being invested in short-term securities returning 1.5 percent annually. What is the NPV of the refund? Use the NPV function in Excel and the PV function in Excel. Answers should be the same.
Answer #1

Assuming that all the costs related to the new issue and refunding the previous issue including premium on call of previous issue, flotation costs of fresh issue and un amortized expenses of previous issue will be allowed to be amortized over the life time of fresh issue.

Summary.

Old Bonds New Bonds
Value of Bonds     90,000,000          90,000,000
Interest rate PA 6% 5.35%
Flotation Costs       1,800,000            1,350,000
Period Yrs 30 22
Period Lapsed Yrs 8 0
Tax Rate 34% 34%
Overlap time One month

To find out the NPV of the New Issue, let us find out the Savings and Expenses the new issue would result in.

Old Bonds New Bonds
Interest rate PA 6% 5.35%
Interest paid PA             5,400,000                4,815,000
Saving in Interest PA                                                      585,000
Un Amortized Flotation costs of old issue             1,320,000                1,320,000
Flotation costs of New Issue                            -                  1,350,000
Total Flotation costs to be amortised in 22 yrs             1,320,000                2,670,000
                  60,000                    121,364
Additional Exp in Flotation costs                                                  61,363.64
Call Premium 0                4,500,000
Amortized for 22 yrs 0              204,545.45
Excess Interest Paid in Overlap time 0                    425,625
Interest Earned in Overlap time 0                    112,500
Net Interest Exp in Overlap time 0                    313,125
One time Expenses of the New Issue (Un amortized)
Net Interest Exp in Overlap time                 313,125
After Tax expense (313125*0.66)           206,662.50
Saving in Interest PA 585000
Net Expenses Increased every yr
Additional Exp in Flotation costs 61363.63
Call premium 204545
Total 265908.63
Net Savings PA 319091.37
After Tax savings (Every Year) 210600.3042

Assuming the Discount rate at 9%.

PV of Net Issue = PV(9%,22,210600.3042)-206662.5 = $ 1,988,577.677- 206662.5 = $ 1,781,915.71

NPV of Net Issue =NPV(9%,J237:J258)+J236 =  $ 1,781,915.71.

Please use the below data for NPV in excel.

-206662.5
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042
210600.3042